A mixture of three gases has a total pressure of 1,380 mmHg at 298 K. The mixture is analyzed and is found to contain 1.27 mol CO2, 3.04 mol CO, and 1.50 mol Ar. What is the partial pressure of Ar?
Answer #1
n = pV/RT = 0.855 atm x 2.5 L/ 0.08206 x 323 K= 0.0807
molecules = 0.0807 x 6.02 x 10^23 =4.86 x 10^22

total moles = 5.81

1.50 / 5.81 = partial pressure Ar/ 1380
partial pressure Ar = 356.3 mm Hg
Answer #2
n = pV/RT = 0.855 atm x 2.5 L/ 0.08206 x 323 K= 0.0807
molecules = 0.0807 x 6.02 x 10^23 =4.86 x 10^22

total moles = 5.81

1.50 / 5.81 = partial pressure Ar/ 1380
partial pressure Ar = 356.3 mm Hg